Please help me with this problem:
In the Olympics, the Canadian team won a total of 22 medals, consisting of gold, silver, and bronze. There were five fewer gold medals than bronze, and 3 fewer bronze than silver medals. How many gold, silver, and bronze medals did Canada win?

Answer by stanbon(48568)

In the Olympics, the Canadian team won a total of 22 medals, consisting of gold, silver, and bronze. There were five fewer gold medals than bronze, and 3 fewer bronze than silver medals. How many gold, silver, and bronze medals did Canada win?
------------------------------
g + s + b = 22
g = b-5
b = s-3
------------------
Rearrange #3:
s = b+3
-----------------
Substitute for s and for g in #1 to get:
(b-5) + (b+3) + b = 22
3b -2 = 22
3b = 24
b = 8 (# of bronze medals)
------------
Then g = b-5 = 8-5 = 3 (# of gold medals)
and s + b+3 = 8+3 = 11 (# of silver medals)
=============================================
Cheers,
Stan H.